# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v3

@@
11email: matteo.ceci@unifi.it 22institutetext: INAF - Osservatorio Astrofisico di Arcetri, Largo E. Fermi 5, I-50125 Firenze, Italy 33institutetext: Scuola Normale Superiore, Piazza dei Cavalieri 7, 56126 Pisa, Italy 44institutetext: Instituto de Radioastronomía y Astrofísica, Universidad Nacional Autónoma de México, Morelia, Michoacán 58089, Mexico 55institutetext: ESO, Karl-Schwarzschild-Str. 2, 85748 Garching bei München, Germany 66institutetext: Instituto de Astrofísica de Canarias, 38205 La Laguna, Tenerife, Spain 77institutetext: Departamento de Astrofísica, Universidad de La Laguna, 38206 La Laguna, Tenerife, Spain 88institutetext: Institute of Physics, GALSPEC laboratory, EPFL, Observatory of Sauverny, Chemin Pegasi 51, 1290 Versoix, Switzerland 99institutetext: AURA for ESA, Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 1010institutetext: INAF, Osservatorio Astronomico di Trieste, Via Tiepolo 11, I-34131 Trieste, Italy 1111institutetext: IFPU - Institute for Fundamental Physics of the Universe, via Beirut 2, I-34151 Trieste, Italy 1212institutetext: University of Trento, Via Sommarive 14, Trento, I-38123, Italy 1313institutetext: Max-Planck-Institut für Extraterrestrische Physik (MPE), Gießenbachstr. 1, D-85748 Garching, Germany 1414institutetext: Observatorio Astronómico Nacional, C/ Alfonso XII 3, 28014 Madrid, Spain 1515institutetext: Dipartimento di Fisica e Astronomia, Alma Mater Studiorum, Università degli Studi di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 1616institutetext: INAF–Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, Via Gobetti 93/3, 40129 Bologna, Italy
MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling
@@
We present a multi-phase analysis of the gas in the circumnuclear region ( 0.90.9 kpc2) of the nearby barred Seyfert 1.8 galaxy NGC 1365, observed as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. Specifically, we combined spatially resolved spectroscopic data from JWST/MIRI, VLT/MUSE and ALMA to provide a multi-phase characterization of the ionized atomic and the warm and cold molecular gas phases. MIRI data enabled the detection of more than 40 mid-IR emission lines from ionized or warm molecular gas. Moment maps show that both cold and warm molecular gas trace the circumnuclear ring, following the rotation of the stellar disk. The ionized gas exhibits flux distributions and kinematics varying depending on the ionization potential (IP). Low-IP species ( 25 eV) mainly trace the rotating disk, while higher-IP species (up to 120 eV) trace the outflowing gas. Both [O III] and [Ne V] trace the nuclear outflow cone towards the southeast. In addition, the [Ne V] line traces the counter-cone of the outflow to the northwest, which is obscured in the optical at these circumnuclear scales and thus undetected in [O III]. Unlike optical diagnostics, spatially resolved mid-IR diagnostics reveal the key role of the active galactic nuclei (AGN) as the source of gas ionization in the central region. We derive the electron density from the [Ne V]/[Ne V] line ratio, finding a median value of (750440) cm-3, consistent with previous estimates obtained from the optical [S II] doublet. Lastly, we apply, for the first time, a fully self-consistent combination of state-of-the-art photoionization and kinematic models (HOMERUN+) to estimate the intrinsic physical outflow properties, kinematics, and energetics - overcoming the limitations of classical methods based on oversimplified assumptions. Exploiting the unprecedented synergy between JWST/MIRI and VLT/MUSE, HOMERUN allows us to simultaneously reproduce the fluxes of over 60 emission lines spanning from the optical to the mid-IR. This unique approach enables us to disentangle the physical conditions of AGN- and star formation-dominated components and robustly estimate the mass of the outflowing gas and other physical properties.
@@
Galactic outflows are known to influence the evolution of their host galaxies by removing gas, enriching the circumngalactic medium, and suppressing star formation in the galaxy (Veilleux et al., 2005; Fabian, 2012; Cresci & Maiolino, 2018; Harrison & Ramos Almeida, 2024). While powerful outflows are commonly observed in luminous active galactic nuclei (AGN) at redshifts (e.g., Cano-Díaz et al., 2012; Fiore et al., 2017), similar signatures are also detected in local, less-powerful AGN host galaxies, where high spatial resolution enabled by their vicinity allows for detailed studies of their physical and kinematic properties (Cresci et al., 2015; Venturi et al., 2017, 2018; Cicone et al., 2018; Mingozzi et al., 2019; Fluetsch et al., 2019; Lutz et al., 2020). These local outflows span multiple gas phases, from ionized atomic to molecular, and show broad, blue-shifted components in emission and/or absorption lines, consistent with being driven by AGN activity.
@@
Ionized outflows, often traced by strong optical emission lines such as [O III], exhibit velocities ranging from several hundreds to a few thousands km s-1. They are composed of warm gas (T = K, ne=cm-3) that expands from the nuclear region out to galactic scales, potentially expelling material from their host galaxy (e.g., Harrison et al., 2014, 2016; Woo et al., 2016; Fiore et al., 2017; Venturi et al., 2018; Förster Schreiber et al., 2019; Kakkad et al., 2020; Venturi et al., 2021; Cresci et al., 2023; Speranza et al., 2024; Marconcini et al., 2025b). However, the impact of ionized winds on star formation (SF) remains debated (e.g., Harrison et al., 2016; Harrison & Ramos Almeida, 2024), as their coupling efficiency with the ambient interstellar medium (ISM) appears to be low. Indeed, the energy and momentum carried by the ionized gas phase are typically sub-dominant when compared to those of other gas phases (e.g. Zubovas & Bourne, 2017; Combes, 2017; Fluetsch et al., 2019, 2021; Mulcahey et al., 2022; Belli et al., 2024). In the context of these findings, the ionized gas component may only trace the larger scale, less massive regions of multi-phase outflows, which are dominated in mass by colder molecular or atomic gas on less extended scales (e.g. Ramos Almeida et al., 2022; Audibert et al., 2023; Venturi et al., 2023).
@@
Cold molecular outflows represent a colder phase of gas (T = K, n cm-3), frequently observed via CO emission lines. Compared to the ionized phase, CO-traced outflows are typically slower, with velocities of a few hundred km s-1, but often carry substantial mass, suggesting they may play a crucial role in regulating SF (Feruglio et al., 2010; Fiore et al., 2017; García-Burillo et al., 2019; Lutz et al., 2020; Fluetsch et al., 2019, 2021). In addition to the cold component, warm molecular outflows (T = K) are typically traced by H2 roto-vibrational transitions in the near- and mid-infrared (IR). H2 emission is particularly prominent in shocked regions, making it a sensitive tracer of warm molecular gas entrained in outflows (Hill & Zakamska, 2014; Richings & Faucher-Giguère, 2018a, b; Riffel et al., 2020; Wright et al., 2023).
@@
NGC 1365, the subject of this work, is a great laboratory to study in detail multi-phase AGN feedback. This dusty barred spiral galaxy (SB(s)b; de Vaucouleurs et al. 1991) is located in the Fornax cluster (Jones & Jones, 1980) at a distance of 19.57 Mpc (Jacobs et al., 2009; Anand et al., 2021) (1″ 95 pc). With a redshift of z = 0.005457 (Bureau et al., 1996), NGC 1365 is classified as a Seyfert 1.8 galaxy (Véron-Cetty & Véron, 2006), with signature of both AGN activity and SF in its central regions.
@@
The AGN-driven ionized outflow in NGC 1365 is observed via extended biconical [O III] emission spanning 2.5 kpc, with line-of-sight velocities up to 170 km s-1 (Venturi et al., 2018). The southeastern (SE) cone, approaching the observer, and the receding northwestern (NW) cone - partially obscured by the galactic disk - are dominated by AGN ionization as traced by diagnostic diagrams (Baldwin et al., 1981; Veilleux & Osterbrock, 1987; Kewley et al., 2006). High-resolution X-ray observations revealed a fast, highly ionized blue-shifted nuclear (spatially unresolved) wind in absorption with velocities of 3000 km s-1, further supporting the scenario of an AGN-driven wind (Risaliti et al., 2005; Braito et al., 2014).
@@
NGC 1365 has a SF rate of 16.9 M⊙ yr-1 (Lee et al., 2022), mainly within a 2 k radius circumnuclear ring identified across optical, radio, and IR wavelengths (Kristen et al., 1997; Forbes & Norris, 1998; Alonso-Herrero et al., 2012). The ring is associated with the inner Lindblad resonance (Lindblad et al., 1996) and shows non-circular bar-driven gas motions (Teuben et al., 1986; Sanchez, 2009), associated with SF-driven ionization (Sharp & Bland-Hawthorn, 2010; Venturi et al., 2018) and large reservoirs of molecular gas ( ; Sakamoto et al., 2007; Gao et al., 2021). Observational campaigns from the PHANGS survey (“Physics at High Angular resolution in Nearby GalaxieS”, Lee et al. 2023) conducted with the James Webb Space Telescope (JWST) and complemented by the Atacama Large Millimeter/submillimeter Array (ALMA; Wootten & Thompson 2009) have revealed that young clusters embedded in the ring drive both ionization of the surrounding medium and localized modifications to the molecular gas phase, where feedback-driven heating and dissociation decrease CO excitation and increase [C I]/CO abundance ratios in their immediate vicinity (Liu et al., 2023; Schinnerer et al., 2023).
@@
Despite hints for the presence of a nuclear radio jet in NGC 1365 (Sandqvist et al., 1995), later studies by Stevens et al. (1999) found no conclusive evidence for its presence. Instead, radio emission is primarily attributed to the elongated star-forming ring rather than a jet.
@@
JWST is transforming our understanding of the outflow properties in the local Universe by providing an unprecedented view of the nuclear dusty regions of AGN (e.g. Hermosa Muñoz et al., 2024a; Perna et al., 2024; Zhang et al., 2024; García-Bernete et al., 2024; Davies et al., 2024; Ulivi et al., 2025; Ceci et al., 2025). This revolution is largely driven by the unique capabilities of the JWST Mid-Infrared Instrument (MIRI; Wright et al., 2015, 2023), including its high spatial resolution across a wide range of wavelengths (from 5 m to 28 m), enabling detailed studies of the energetics and kinematics of gas flows in obscured nuclear regions of AGN. The wide IR spectral range of JWST spanning many atomic and molecular gas transitions, combined with its spatial and spectral resolution, provides essential insights into the multi-phase circumnuclear medium. These unprecedented capabilities enable multi-phase, co-spatial analyses when combined with complementary data from ground-based facilities, such as the Multi Unit Spectroscopic Explorer (MUSE) at the ESO Very Large Telescope (VLT; Bacon et al. 2010) and ALMA, transforming our understanding of feedback mechanisms in local AGN.
@@
For NGC 1365, we combine data from the Medium-Resolution Spectrometer (MRS; Rieke et al. 2015; Labiano et al. 2021) of MIRI with integral-field optical spectroscopy from VLT/MUSE and millimeter observations from ALMA. This multi-wavelength approach enables us to trace the multi-phase gas across a wide range of ionization states and densities. In particular, it allows for a detailed characterization of the kinematics, morphology, and energetics of the circumnuclear environment in NGC 1365, ultimately contributing to a deeper understanding of AGN feedback and its impact on galaxy evolution.
@@ 2 Overview of observations and data analysis
NGC 1365 was observed on 8th December 2024 as part of the “Mid-IR Activity of Circumnuclear Line Emission” (MIRACLE; JWST GO program 6138; Co-PIs: C. Marconcini and A. Feltre), aimed at tracing the mid-IR emission by exploiting MIRI/MRS data of the circumnuclear region of local AGN in the 5-28 m wavelength range. The MRS mode of MIRI covers a total wavelength range of 4.9–27.9 m, divided into four integral field units (IFUs), also referred to as channels (Ch1, Ch2, Ch3, and Ch4, hereafter), each further subdivided into three bands (SHORT, MEDIUM, and LONG). These channels cover slightly different fields of view (FoVs), from 3.2″ 3.7″ in Ch1 to 6.6″ 7.7″ in Ch4, at varying pixel size (from 0.13″ in Ch1 to 0.35″ in Ch4) and resolving powers (from 3700 to 1500; see e.g. Labiano et al., 2021; Argyriou et al., 2023). As shown in Law et al. (2023), the average FWHM of the MIRI/MRS PSF ranges from 0.4″ in Ch1 to 0.9″ in Ch4. We summarize the data reduction in Appendix A.1, referring to Marconcini et al. (2025a) for a detailed description. As a result of this process, the pipeline produced 12 datacubes, one for each sub-band.
@@
The science observations targeted the nuclear region of NGC 1365 but were deliberately offset to include the inner part of the SE approaching outflow. As shown in the lower panels of Fig. 1, the different MIRI FoVs of each channel are not exactly centered on the nucleus, but shifted toward the SE to better capture the outflow region. This choice was motivated by the primary goal of our MIRACLE project of characterizing multi-phase gas outflows and nuclear properties. In the lower right panel of Fig. 1, we show the nuclear region of NGC 1365 in the MIRI FoV. We extracted integrated spectra from the 0.7″ radius circular apertures, and we present them in Fig. 2. The green (blue) spectrum is extracted from the circular region shown in Fig. 1 where the outflow (stellar disk) emission is dominant, as is shown by the flux maps in Fig. 3. Note that the stellar disk spectrum is contaminated by outflow emission, as indicated by the presence of high-excitation lines. We detected more than 40 emission lines, with a signal-to-noise ratio (S/N) 5, in the spectral range covered by MIRI, mostly tracing ionized gas, with an ionization potential IP - defined as the energy required to create the relevant species - ranging from a few eV to more than 100 eV (see Table LABEL:tab:_table_list_emission_lines). Additionally, we also detected two hydrogen recombination lines, namely Pf and Hu (i.e. H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) and seven H2 pure-rotational transitions, from 0-0 S(7) to 0-0 S(1).
@@
As shown in Fig. 2, the MIRI spectra also exhibit at least 15 bright PAH features in the wavelength range between 5 - 18 m, which we will investigate in a dedicated forthcoming paper. To obtain these spectra, for the first time we combined the emission from all 12 MIRI datacubes, taking into account the different pixel sizes and ensuring a proper flux conservation among different bands by applying a scaling factor to align the flux levels between adjacent bands. A detailed description of this novel procedure is presented in the Appendix B. This algorithm, written in Python, is available for download111https://github.com/matteo-ceci/JWST-MIRI_stitching_code.
@@
In Appendix C.1, we analyzed the MIRI MRS datacubes using a customized python script to enhance the S/N with spatial smoothing, subtract the local continuum around all the brightest emission lines and then fit them spaxel by spaxel. The fitted emission lines are listed in Table LABEL:tab:_table_list_emission_lines.
@@
The MUSE IFU observations of NGC 1365 were obtained on the 12th October 2014, under program 094.B-0321(A) (PI A. Marconi) and are part of the “Measuring AGN Under MUSE” (MAGNUM) survey (Cresci et al., 2015; Venturi et al., 2017, 2018; Mingozzi et al., 2019; Marconcini et al., 2023, 2025b). We applied the same data reduction pipeline that has been extensively applied in previous works (e.g. Venturi et al., 2018; Mingozzi et al., 2019), and which has been validated and described in detail therein. To trace the cold molecular gas, we used archival ALMA 12-m Band 7 observations (program 2016.1.00296.S, PI F.Combes; Combes et al. 2019) of the CO J= transition at 345.796GHz (rest-frame). The observation description, data reduction and emission line fitting procedure of the MUSE and ALMA data can be found in Appendices A.2-C.2 and A.3-C.3, respectively.

### v4

# Key Words.: galaxies: Seyfert - galaxies: ISM - galaxies: active - ISM: kinematics and dynamics - ISM: jets and outflows1 Introduction

@@
11email: matteo.ceci@unifi.it 22institutetext: INAF - Osservatorio Astrofisico di Arcetri, Largo E. Fermi 5, I-50125 Firenze, Italy 33institutetext: European Southern Observatory, Karl-Schwarzschild Straße 2, D-85748 Garching bei München, Germany 44institutetext: Scuola Normale Superiore, Piazza dei Cavalieri 7, 56126 Pisa, Italy 55institutetext: Instituto de Radioastronomía y Astrofísica, Universidad Nacional Autónoma de México, Morelia, Michoacán 58089, Mexico 66institutetext: Instituto de Astrofísica de Canarias, 38205 La Laguna, Tenerife, Spain 77institutetext: Departamento de Astrofísica, Universidad de La Laguna, 38206 La Laguna, Tenerife, Spain 88institutetext: Institute of Physics, GALSPEC laboratory, EPFL, Observatory of Sauverny, Chemin Pegasi 51, 1290 Versoix, Switzerland 99institutetext: AURA for ESA, Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 1010institutetext: INAF, Osservatorio Astronomico di Trieste, Via Tiepolo 11, I-34131 Trieste, Italy 1111institutetext: IFPU - Institute for Fundamental Physics of the Universe, via Beirut 2, I-34151 Trieste, Italy 1212institutetext: University of Trento, Via Sommarive 14, Trento, I-38123, Italy 1313institutetext: Max-Planck-Institut für Extraterrestrische Physik (MPE), Gießenbachstr. 1, D-85748 Garching, Germany 1414institutetext: Observatorio Astronómico Nacional, C/ Alfonso XII 3, 28014 Madrid, Spain 1515institutetext: Dipartimento di Fisica e Astronomia, Alma Mater Studiorum, Università degli Studi di Bologna, Via Gobetti 93/2, 40129 Bologna, Italy 1616institutetext: INAF–Osservatorio di Astrofisica e Scienza dello Spazio di Bologna, Via Gobetti 93/3, 40129 Bologna, Italy
MIRACLE
@@
We present a multiphase analysis of the gas in the circumnuclear region ( 0.90.9 kpc2) of the nearby barred Seyfert 1.8 galaxy NGC 1365, observed as part of the Mid-IR Activity of Circumnuclear Line Emission (MIRACLE) program. Specifically, we combined spatially resolved spectroscopic data from JWST/MIRI, VLT/MUSE, and ALMA to provide a multiphase characterization of the ionized atomic and the warm and cold molecular gas phases. MIRI data enabled the detection of more than 40 mid-IR emission lines from ionized or warm molecular gas. Moment maps show that both cold and warm molecular gas trace the circumnuclear ring, following the rotation of the stellar disk. The ionized gas exhibits flux distributions and kinematics that vary depending on the ionization potential (IP). Low-IP species ( 25 eV) mainly trace the rotating disk, while higher-IP species (up to 120 eV) trace the outflowing gas. Both [O III] and [Ne V] trace the nuclear outflow cone toward the southeast. In addition, the [Ne V] line traces the counter-cone of the outflow to the northwest, which is obscured in the optical at these circumnuclear scales, and is thus undetected in [O III]. Unlike optical diagnostics, spatially resolved mid-IR diagnostics reveal the key role of the active galactic nucleus (AGN) as the source of gas ionization in the central region. We derived the electron density from the [Ne V]/[Ne V] line ratio, finding a median value of (750440) cm-3, consistent with previous estimates obtained from the optical [S II] doublet. Lastly, we applied, for the first time, a fully self-consistent combination of state-of-the-art photoionization and kinematic models (HOMERUN+) to estimate the intrinsic physical outflow properties, kinematics, and energetics - overcoming the limitations of classical methods based on oversimplified assumptions. Exploiting the unprecedented synergy between JWST/MIRI and VLT/MUSE, HOMERUN allows us to simultaneously reproduce the fluxes of over 60 emission lines spanning from the optical to the mid-IR. This unique approach enables us to disentangle the physical conditions of AGN- and star formation-dominated components and robustly estimate the mass of the outflowing gas and other physical properties.
@@
Galactic outflows are known to influence the evolution of their host galaxies by removing gas, enriching the circumngalactic medium, and suppressing star formation (SF) in the galaxy (Veilleux et al., 2005; Fabian, 2012; Cresci and Maiolino, 2018; Harrison and Ramos Almeida, 2024). While powerful outflows are commonly observed in luminous active galactic nuclei (AGNs) at redshifts (e.g., Cano-Díaz et al., 2012; Fiore et al., 2017), similar signatures are also detected in local, less powerful AGN-hosted galaxies, where the high spatial resolution enabled by their vicinity allows for detailed studies of their physical and kinematic properties (Cresci et al., 2015; Venturi et al., 2017, 2018; Cicone et al., 2018; Mingozzi et al., 2019; Fluetsch et al., 2019; Lutz et al., 2020). These local outflows span multiple gas phases, from ionized atomic to molecular, and show broad, blueshifted components in emission and/or absorption lines, consistent with being driven by AGN activity.
@@
Ionized outflows, often traced by strong optical emission lines such as [O III], exhibit velocities ranging from several hundred to a few thousand kilometers per second. They are composed of warm gas (T = K, ne=cm-3) that expands from the nuclear region out to galactic scales, potentially expelling material from their host galaxy (e.g., Harrison et al., 2014, 2016; Woo et al., 2016; Fiore et al., 2017; Venturi et al., 2018; Förster Schreiber et al., 2019; Kakkad et al., 2020; Venturi et al., 2021; Cresci et al., 2023; Speranza et al., 2024; Marconcini et al., 2025b). However, the impact of ionized winds on SF remains debated (e.g., Harrison et al., 2016; Harrison and Ramos Almeida, 2024), as their coupling efficiency with the ambient interstellar medium (ISM) appears to be low. Indeed, the energy and momentum carried by the ionized gas phase are typically subdominant when compared to those of other gas phases (e.g., Zubovas and Bourne, 2017; Combes, 2017; Fluetsch et al., 2019, 2021; Mulcahey et al., 2022; Belli et al., 2024). In the context of these findings, the ionized gas component may only trace the larger-scale, less massive regions of multiphase outflows, which are dominated in mass by colder molecular or atomic gas on less extended scales (e.g., Ramos Almeida et al., 2022; Audibert et al., 2023; Venturi et al., 2023).
@@
Cold molecular outflows represent a colder phase of gas (T = K, n cm-3), frequently observed via CO emission lines. Compared to the ionized phase, CO-traced outflows are typically slower, with velocities of a few hundred kilometers per second, but often carry substantial mass, suggesting they may play a crucial role in regulating SF (Feruglio et al., 2010; Fiore et al., 2017; García-Burillo et al., 2019; Lutz et al., 2020; Fluetsch et al., 2019, 2021). In addition to the cold component, warm molecular outflows (T = K) are typically traced by H2 roto-vibrational transitions in the near- and mid-infrared (IR). H2 emission is particularly prominent in shocked regions, making it a sensitive tracer of warm molecular gas entrained in outflows (Hill and Zakamska, 2014; Richings and Faucher-Giguère, 2018a, b; Riffel et al., 2020; Wright et al., 2023).
@@
NGC 1365, the subject of this work, is a great laboratory in which to study in detail multiphase AGN feedback. This dusty barred spiral galaxy (SB(s)b; de Vaucouleurs et al. 1991) is located in the Fornax cluster (Jones and Jones, 1980) at a distance of 19.57 Mpc (Jacobs et al., 2009; Anand et al., 2021) (1″ 95 pc). With a redshift of z = 0.005457 (Bureau et al., 1996), NGC 1365 is classified as a Seyfert 1.8 galaxy (Véron-Cetty and Véron, 2006), with signatures of both AGN activity and SF in its central regions.
@@
The AGN-driven ionized outflow in NGC 1365 is observed via extended biconical [O III] emission spanning 2.5 kpc, with line-of-sight velocities of up to 170 km s-1 (Venturi et al., 2018). The southeastern (SE) cone, approaching the observer, and the receding northwestern (NW) cone – partially obscured by the galactic disk – are dominated by AGN ionization as traced by diagnostic diagrams (Baldwin et al., 1981; Veilleux and Osterbrock, 1987; Kewley et al., 2006). High-resolution X-ray observations revealed a fast, highly ionized blueshifted nuclear (spatially unresolved) wind in absorption with velocities of 3000 km s-1, further supporting the scenario of an AGN-driven wind (Risaliti et al., 2005; Braito et al., 2014).
@@
NGC 1365 has a SF rate of 16.9 M⊙ yr-1 (Lee et al., 2022), mainly within a 2 kpc radius circumnuclear ring identified across optical, radio, and IR wavelengths (Kristen et al., 1997; Forbes and Norris, 1998; Alonso-Herrero et al., 2012). The ring is associated with the inner Lindblad resonance (Lindblad et al., 1996) and shows noncircular bar-driven gas motions (Teuben et al., 1986; Sanchez, 2009), associated with SF-driven ionization (Sharp and Bland-Hawthorn, 2010; Venturi et al., 2018) and large reservoirs of molecular gas ( ; Sakamoto et al., 2007; Gao et al., 2021). Observational campaigns from the PHANGS survey (“Physics at High Angular resolution in Nearby GalaxieS”, Lee et al. 2023) conducted with the James Webb Space Telescope (JWST) and complemented by the Atacama Large Millimeter/submillimeter Array (ALMA; Wootten and Thompson 2009) have revealed that young clusters embedded in the ring drive both ionization of the surrounding medium and localized modifications to the molecular gas phase, where feedback-driven heating and dissociation decrease CO excitation and increase [C I]/CO abundance ratios in their immediate vicinity (Liu et al., 2023; Schinnerer et al., 2023).
@@
Despite hints of a nuclear radio jet in NGC 1365 (Sandqvist et al., 1995), later studies by Stevens et al. (1999) found no conclusive evidence of its presence. Instead, radio emission is primarily attributed to the elongated star-forming ring rather than a jet.
@@
JWST is transforming our understanding of the outflow properties in the local Universe by providing an unprecedented view of the nuclear dusty regions of AGNs (e.g., Hermosa Muñoz et al., 2025; Perna et al., 2024; Zhang et al., 2024; García-Bernete et al., 2024; Davies et al., 2024; Ulivi et al., 2025; Ceci et al., 2025). This revolution is largely driven by the unique capabilities of the JWST Mid-Infrared Instrument (MIRI; Wright et al., 2015, 2023), including its high spatial resolution across a wide range of wavelengths (from 5 m to 28 m), enabling detailed studies of the energetics and kinematics of gas flows in obscured nuclear regions of AGNs. The wide IR spectral range of JWST spanning many atomic and molecular gas transitions, combined with its spatial and spectral resolution, provides essential insights into the multiphase circumnuclear medium. These unprecedented capabilities enable multiphase, co-spatial analyses when combined with complementary data from ground-based facilities, such as the Multi Unit Spectroscopic Explorer (MUSE) at the ESO Very Large Telescope (VLT; Bacon et al. 2010) and ALMA, transforming our understanding of feedback mechanisms in local AGNs.
@@
For NGC 1365, we combined data from the Medium-Resolution Spectrometer (MRS; Rieke et al. 2015; Labiano et al. 2021) of MIRI with integral-field optical spectroscopy from VLT/MUSE and millimeter observations from ALMA. This multiwavelength approach enables us to trace the multiphase gas across a wide range of ionization states and densities. In particular, it allows for a detailed characterization of the kinematics, morphology, and energetics of the circumnuclear environment in NGC 1365, ultimately contributing to a deeper understanding of AGN feedback and its impact on galaxy evolution.
@@ 2 Overview of observations and data analysis
NGC 1365 was observed on 8 December 2024 as part of the “Mid-IR Activity of Circumnuclear Line Emission” (MIRACLE; JWST GO program 6138; Co-PIs: C. Marconcini and A. Feltre), aimed at tracing the mid-IR emission by exploiting MIRI/MRS data of the circumnuclear region of local AGNs in the 5-28 m wavelength range. The MRS mode of MIRI covers a total wavelength range of 4.9–27.9 m, divided into four integral field units (IFUs), also referred to as channels (Ch1, Ch2, Ch3, and Ch4, hereafter), each further subdivided into three bands (SHORT, MEDIUM, and LONG). These channels cover slightly different fields of view (FoVs), from 3.2″ 3.7″ in Ch1 to 6.6″ 7.7″ in Ch4, at varying pixel sizes (from 0.13″ in Ch1 to 0.35″ in Ch4) and resolving powers (from 3700 to 1500; see e.g., Labiano et al., 2021; Argyriou et al., 2023). As shown in Law et al. (2023), the average FWHM of the MIRI/MRS PSF ranges from 0.4″ in Ch1 to 0.9″ in Ch4. We summarize the data reduction in Appendix A.1, referring to Marconcini et al. (2025a) for a detailed description. As a result of this process, the pipeline produced 12 datacubes, one for each sub-band.
@@
The science observations targeted the nuclear region of NGC 1365 but were deliberately offset to include the inner part of the SE approaching outflow. As is shown in the lower panels of Fig. 1, the different MIRI FoVs of each channel are not exactly centered on the nucleus, but shifted toward the SE to better capture the outflow region. This choice was motivated by the primary goal of our MIRACLE project of characterizing multiphase gas outflows and nuclear properties. In the lower right panel of Fig. 1, we show the nuclear region of NGC 1365 in the MIRI FoV. We extracted integrated spectra from the 0.7″ radius circular apertures, and we present them in Fig. 2. The green (blue) spectrum is extracted from the circular region shown in Fig. 1 where the outflow (stellar disk) emission is dominant, as is shown by the flux maps in Fig. 3. Note that the stellar disk spectrum is contaminated by the outflow emission, as is indicated by the presence of high-excitation lines. We detected more than 40 emission lines, with a signal-to-noise ratio (S/N) 5, in the spectral range covered by MIRI, mostly tracing ionized gas, with an ionization potential (IP) – defined as the energy required to create the relevant species – ranging from a few electronvolts to more than 100 eV (see Table LABEL:tab:_table_list_emission_lines). Additionally, we also detected two hydrogen recombination lines; namely, Pf and Hu (i.e., H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) and seven H2 pure-rotational transitions, from 0-0 S(7) to 0-0 S(1).
@@
As is shown in Fig. 2, the MIRI spectra also exhibit at least 15 bright PAH features in the wavelength range between 5 - 18 m, which we will investigate in a dedicated forthcoming paper. To obtain these spectra, for the first time we combined the emission from all 12 MIRI datacubes, taking into account the different pixel sizes and ensuring a proper flux conservation among different bands by applying a scaling factor to align the flux levels between adjacent bands. A detailed description of this novel procedure is presented in the Appendix B. This algorithm, written in Python, is available for download111https://github.com/matteo-ceci/JWST-MIRI_stitching_code.
@@
In Appendix C.1, we analyzed the MIRI MRS datacubes using a customized python script to enhance the S/N with spatial smoothing, subtract the local continuum around all the brightest emission lines and then fit them spaxel by spaxel. The fit emission lines are listed in Table LABEL:tab:_table_list_emission_lines.
@@
The MUSE IFU observations of NGC 1365 were obtained on the 12th October 2014, under program 094.B-0321(A) (PI A. Marconi) and are part of the “Measuring AGN Under MUSE” (MAGNUM) survey (Cresci et al., 2015; Venturi et al., 2017, 2018; Mingozzi et al., 2019; Marconcini et al., 2023, 2025b). We applied the same data reduction pipeline that has been extensively applied in previous works (e.g., Venturi et al., 2018; Mingozzi et al., 2019), and which has been validated and described in detail therein. To trace the cold molecular gas, we used archival ALMA 12-m Band 7 observations (program 2016.1.00296.S, PI F.Combes; Combes et al. 2019) of the CO J= transition at 345.796GHz (rest-frame). The observation description, data reduction and emission line fitting procedure of the MUSE and ALMA data can be found in Appendices A.2-C.2 and A.3-C.3, respectively.